# PoisonedFL: Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v1

@@ Appendix A Additional Experimental Results
Figure 10. The testing error rate of a global model as a function of the magnitude (measured by $\ell_{2}$ norm) of the noise added to it along a random update direction. We assume the initial global model is accurate and trained under no attack. The dataset is Purchase and FL defense is Median.
Figure 11. The testing error rate of a global model as a function of the predefined norm $b$ used to rescale the total aggregated model update. The dataset is Purchase and FL defense is Median.
A.1. Analyzing the Effectiveness of PoisonedFL
Table 4. The fraction (%) of dimensions of the aggregated model update whose signs match with the sign vector $\bm{s}$ under PoisonedFL. The dataset is Purchase.
@@ A.1. Analyzing the Effectiveness of PoisonedFL
Table 5. $\ell_{2}$ norm (×$\times 10^{2}$) of the total aggregated model updates $\left\|\sum_{t^{\prime}=1}^{t}\bm{g}^{t^{\prime}}\right\|$ (i.e., $\left\|\bm{w}^{t}-\bm{w}^{0}\right\|$) for different attacks and defenses on the Purchase dataset.
@@
Table 6. The fraction (mean $\pm$ standard deviation, %) of malicious model updates in PoisonedFL that are not filtered out by FLAME and Multi-Krum (i.e., false negative rate) across rounds. The dataset is Purchase.
@@
Table 7. The normalized trust score (mean $\pm$ standard deviation) of the malicious/benign model updates in FLTrust against PoisonedFL across rounds. The positive normalized trust scores indicate that the malicious model updates are not entirely filtered out. The dataset is Purchase.
@@
A.2. Impact of the Sign Vector
Table 8. Testing error rate of the global model under PoisonedFL with various random sign vectors obtained from different random seeds on CIFAR-10.
@@ Appendix B Additional Details of Experimental Setup
B.1. Datasets
We use the following five datasets from different domains.
MNIST (LeCun et al., 1998):  MNIST is a 10-class handwritten digits classification dataset. It comprises 60,000 training examples and 10,000 testing examples. In FL, the training data is typically not independently and identically distributed (non-IID) among clients. Following (Cao and Gong, 2022; Fang et al., 2020), we distribute the training examples among clients based on a non-IID degree of $q=0.5$ by default. We train a convolutional neural network (CNN) for MNIST dataset. The architecture of CNN is shown in Table 9 in Appendix.
FashionMNIST (Xiao et al., 2017):  FashionMNIST is a benchmark dataset of Zalando’s article images, which contains 60,000 training examples and 10,000 testing examples. For FashionMNIST, the degree of non-IID is also set to $q=0.5$ by default. We use the same CNN architecture for FashionMNIST dataset as employed in MNIST dataset.
Purchase (pur, 2021):  Purchase is a 100-class customer purchase style prediction dataset. Each input consists of 600 binary features. Following (Cao and Gong, 2022), we split the total 197,324 examples into 180,000 training examples and 17,324 testing examples. Following (Cao and Gong, 2022), we evenly distribute the training examples to clients and train a fully connected neural network.
CIFAR-10 (Krizhevsky et al., 2009):  CIFAR-10 is 10-class color image classification dataset. It comprises 50,000 training examples and 10,000 testing examples. To consider a different degree of non-IID, we set $q=0.2$ for CIFAR-10 dataset. We train a CNN for CIFAR-10, whose architecture is shown in Table 10.
FEMNIST (Caldas et al., 2018):  FEMNIST is a 62-class classification dataset. The dataset is already distributed among 3,550 clients with a total of 805,263 examples. We randomly sample 1,200 clients. We train a CNN for FEMNIST, whose architecture is shown in Table 11.
B.2. FL Settings
@@
We primarily establish the FL setting in a cross-device scenario, which is more feasible for model poisoning attacks, as discussed in Section 8. Therefore, we assume $1,200$ genuine clients and $20\%$ fake clients. The learning rates for the five datasets are selected within the range of $0.01$ to $0.1$ for optimal training effectiveness. Specifically, the learning rates for MNIST, FashionMNIST, and FEMNIST are set to $0.01$, while CIFAR-10 is set to $0.03$, and Purchase is set to $0.1$. The number of training rounds is configured to ensure full convergence under various scenarios: $6,000$ for MNIST, FashionMNIST, and FEMNIST, and $10,000$ for CIFAR-10 and FEMNIST. The batch size is set to $32$ for MNIST, FashionMNIST, CIFAR-10, and FEMNIST, and to $128$ for Purchase, based on the number of local data samples.
B.3. Defenses
@@
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules (Multi-Krum (Blanchard et al., 2017), Median (Yin et al., 2018), Trimmed Mean (Yin et al., 2018), Norm Bound (Sun et al., 2019), FLTrust (Cao et al., 2021a), and FLAME (Nguyen et al., 2022)), one provably robust defense (FLCert (Cao et al., 2022)) and one malicious clients detection method (FLDetector (Zhang et al., 2022)). We also consider the non-robust FedAvg (McMahan et al., 2017) as a baseline.
@@
Multi-Krum (Blanchard et al., 2017):  Multi-Krum uses an iterative method to select a subset of clients’ model updates. In each step, it selects the model update that has the smallest sum of Euclidean distance to its $n-2$ neighbors, where $n$ is the number of genuine clients. This selection process continues until $n$ clients are chosen. Finally, the server computes the average of the $n$ selected model updates as the aggregated model update.
@@
Median (Yin et al., 2018):  In Median aggregation rule, the server computes the coordinate-wise median of the clients’ model updates as the aggregated model update.
@@
Trimmed Mean (TrMean) (Yin et al., 2018):  For each dimension, the server first removes the largest $m$ values and smallest $m$ values, and then computes the average of the remaining $n-m$ values, where $m$ is the number of fake clients.
@@
Norm Bound (Sun et al., 2019):  In Norm Bound aggregation rule, the server first clips each client’s model update to have a predetermined norm and then computes the average of the clipped model updates as the aggregated model update. We follow prior work (Shejwalkar et al., 2022) to set the predetermined norm as the average norm of the genuine model updates in each training round.
@@
FLTrust (Cao et al., 2021a):  FLTrust assumes that the server has a small, clean dataset to bootstrap trust. We assume the server’s dataset includes training examples selected uniformly at random. The size of the server’s dataset is 100 for MNIST, FashionMNIST, and CIFAR-10, and 200 for Purchase and FEMNIST.
@@
FLAME (Nguyen et al., 2022):  In each training round, FLAME considers the cosine similarity between clients’ local models (i.e., a client’s model update + current global model) to divide clients into clusters and filter out the clusters containing potentially malicious model updates.
@@
FLCert (Cao et al., 2022):  FLCert divides clients into multiple groups and trains a global model for each group using any existing FL algorithm (we use Median (Yin et al., 2018)). Given a testing input, FLCert predicts its label based on the majority vote among these global models. We divide the clients into 10 disjoint groups uniformly at random in our experiments.
@@
FLDetector (Zhang et al., 2022):  FLDetector aims to detect malicious clients during training. We first apply FLDetector with the full participation of clients for detection for 300 communication rounds. Following (Zhang et al., 2022), if some clients are detected as malicious, we remove them from the system. Then we re-train a global model from scratch based on the remaining clients using Median (Yin et al., 2018) aggregation rule with the default setting.
Table 9. Model architecture for MNIST and FashionMNIST.
@@
Table 10. Model architecture for CIFAR-10.
@@
Table 11. Model architecture for FEMNIST.
@@ B.4. Compared Attacks
We compare our PoisonedFL with seven attacks, including five attacks (Fang (Fang et al., 2020), Opt. Fang (Shejwalkar and Houmansadr, 2021), LIE (Baruch et al., 2019), Min-Max (Shejwalkar and Houmansadr, 2021), and Min-Sum (Shejwalkar and Houmansadr, 2021)) that require genuine clients’ information and two attacks (Random (Cao and Gong, 2022) and MPAF (Cao and Gong, 2022)) that do not require such information. Note that when applying the first category of attacks to craft the malicious model updates on fake clients, we assume that the attacker has access to the model updates of all genuine clients and the aggregation rule used by the server. In Section 8, we also study the scenarios where an attacker uses the global models to reconstruct synthetic data on the fake clients and uses the local models trained on them to perform attacks. We do not make these assumptions for Random, MPAF, and our PoisonedFL. Therefore, we give advantages to the first category of attacks.
@@
Fang (Fang et al., 2020):  In this attack, the attacker crafts malicious model updates for fake clients such that the aggregated model update after attack deviates substantially from the before-attack one. Fang has different versions for different aggregation rules. We apply their Krum attack for Multi-Krum defense and TrMean attack for others.
@@
Opt. Fang (Shejwalkar and Houmansadr, 2021):  Following Fang, this attack also aims to maximize the deviation between the after-attack aggregated model update and before-attack one, but it uses a different way to solve the malicious model update. Specifically, a malicious model update is a variant of the average of all genuine clients’ model updates.
@@
A little is enough (LIE) (Baruch et al., 2019):  In LIE attack, fake clients craft their malicious model updates by adding a small amount of noise to the average of genuine model updates. Specifically, for each dimension $j$, LIE calculates the mean $\mu_{j}$ and standard deviation $\sigma_{j}$ among the genuine model updates, and the dimension $j$ of the malicious model update is set as $\mu_{j}+0.74\cdot\sigma_{j}$.
@@
Min-Max (Shejwalkar and Houmansadr, 2021):  In this attack, a fake client crafts its malicious model update such that its distance to any genuine model update is no larger than the maximum distance between any two genuine model updates.
@@
Min-Sum (Shejwalkar and Houmansadr, 2021):  Each fake client crafts its malicious model update such that the sum of the distances between the malicious model update and all genuine model updates is not larger than the sum of distances between any genuine model update and other genuine model updates.
@@
Random (Cao and Gong, 2022):  In this attack, each fake client $i\in[n+1,n+m]$ sends a scaled random vector $\bm{g}_{i}^{t}=\lambda\cdot\bm{\epsilon}$ to the server, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\lambda$ is a scaling factor. Following prior work (Cao and Gong, 2022), we set $\lambda=1e6$ in our experiments.
@@
MPAF (Cao and Gong, 2022):  In MPAF attack, during training round $t$, fake client $i\in[n+1,n+m]$ crafts its malicious model update as $\bm{g}_{i}^{t}=\lambda\cdot(\bm{w}^{\prime}-\bm{w}^{t-1}$), where $\bm{w}^{\prime}$ is an attacker-chosen random model, $\bm{w}^{t-1}$ is the global model in the previous training round $t-1$, and $\lambda$ is a scaling factor. Following prior work (Cao and Gong, 2022), we set $\lambda=1e6$ in our experiments.

### v2

@@ Appendix A Additional Experimental Results
Figure 10: The testing error rate of a global model as a function of the magnitude (measured by $\ell_{2}$ norm) of the noise added to it along a random update direction. We assume the initial global model is accurate and trained under no attack. The dataset is Purchase and FL defense is Median.
Figure 11: The testing error rate of a global model as a function of the predefined norm $b$ used to rescale the total aggregated model update. The dataset is Purchase and FL defense is Median.
A.1 Analyzing the Effectiveness of PoisonedFL
Table 4: The fraction (%) of dimensions of the aggregated model update whose signs match with the sign vector $\bm{s}$ under PoisonedFL. The dataset is Purchase.
@@ A.1 Analyzing the Effectiveness of PoisonedFL
Table 5: $\ell_{2}$ norm (×$\times 10^{2}$) of the total aggregated model updates $\left\|\sum_{t^{\prime}=1}^{t}\bm{g}^{t^{\prime}}\right\|$ (i.e., $\left\|\bm{w}^{t}-\bm{w}^{0}\right\|$) for different attacks and defenses on the Purchase dataset.
@@
Table 6: The fraction (mean $\pm$ standard deviation, %) of malicious model updates in PoisonedFL that are not filtered out by FLAME and Multi-Krum (i.e., false negative rate) across rounds. The dataset is Purchase.
@@
Table 7: The normalized trust score (mean $\pm$ standard deviation) of the malicious/benign model updates in FLTrust against PoisonedFL across rounds. The positive normalized trust scores indicate that the malicious model updates are not entirely filtered out. The dataset is Purchase.
@@
A.2 Impact of the Sign Vector
Table 8: Testing error rate of the global model under PoisonedFL with various random sign vectors obtained from different random seeds on CIFAR-10.
@@ Appendix B Additional Details of Experimental Setup
B.1 FL Settings
@@
We primarily establish the FL setting in a cross-device scenario, which is more feasible for model poisoning attacks, as discussed in Section 8. Therefore, we assume $1,200$ genuine clients and $20\%$ fake clients. The learning rates for the five datasets are selected within the range of $0.01$ to $0.1$ for optimal training effectiveness. Specifically, the learning rates for MNIST, FashionMNIST, and FEMNIST are set to $0.01$, while CIFAR-10 is set to $0.03$, and Purchase is set to $0.1$. The number of training rounds is configured to ensure full convergence under various scenarios: $6,000$ for MNIST, FashionMNIST, and FEMNIST, and $10,000$ for CIFAR-10 and FEMNIST. The batch size is set to $32$ for MNIST, FashionMNIST, CIFAR-10, and FEMNIST, and to $128$ for Purchase, based on the number of local data samples. We conduct our experiments using 16 GeForce RTX 3090 GPUs.
B.2 Defenses
@@
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules (Multi-Krum [12], Median [17], Trimmed Mean [17], Norm Bound [16], FLTrust [13], and FLAME [15]), one provably robust defense (FLCert [14]) and one malicious clients detection method (FLDetector [18]). We also consider the non-robust FedAvg [2] as a baseline.
@@
Multi-Krum [12]:  Multi-Krum uses an iterative method to select a subset of clients’ model updates. In each step, it selects the model update that has the smallest sum of Euclidean distance to its $n-2$ neighbors, where $n$ is the number of genuine clients. This selection process continues until $n$ clients are chosen. Finally, the server computes the average of the $n$ selected model updates as the aggregated model update.
@@
Median [17]:  In Median aggregation rule, the server computes the coordinate-wise median of the clients’ model updates as the aggregated model update.
@@
Trimmed Mean (TrMean) [17]:  For each dimension, the server first removes the largest $m$ values and smallest $m$ values, and then computes the average of the remaining $n-m$ values, where $m$ is the number of fake clients.
@@
Norm Bound [16]:  In Norm Bound aggregation rule, the server first clips each client’s model update to have a predetermined norm and then computes the average of the clipped model updates as the aggregated model update. We follow prior work [20] to set the predetermined norm as the average norm of the genuine model updates in each training round.
@@
FLTrust [13]:  FLTrust assumes that the server has a small, clean dataset to bootstrap trust. We assume the server’s dataset includes training examples selected uniformly at random. The size of the server’s dataset is 100 for MNIST, FashionMNIST, and CIFAR-10, and 200 for Purchase and FEMNIST.
@@
FLAME [15]:  In each training round, FLAME considers the cosine similarity between clients’ local models (i.e., a client’s model update + current global model) to divide clients into clusters and filter out the clusters containing potentially malicious model updates.
@@
FLCert [14]:  FLCert divides clients into multiple groups and trains a global model for each group using any existing FL algorithm (we use Median [17]). Given a testing input, FLCert predicts its label based on the majority vote among these global models. We divide the clients into 10 disjoint groups uniformly at random in our experiments.
@@
FLDetector [18]:  FLDetector aims to detect malicious clients during training. We first apply FLDetector with the full participation of clients for detection for 300 communication rounds. Following [18], if some clients are detected as malicious, we remove them from the system. Then we re-train a global model from scratch based on the remaining clients using Median [17] aggregation rule with the default setting.
Table 9: Model architecture for MNIST and FashionMNIST.
@@
Table 10: Model architecture for CIFAR-10.
@@
Table 11: Model architecture for FEMNIST.
@@ B.3 Compared Attacks
We compare our PoisonedFL with seven attacks, including five attacks (Fang [10], Opt. Fang [11], LIE [8], Min-Max [11], and Min-Sum [11]) that require genuine clients’ information and two attacks (Random [9] and MPAF [9]) that do not require such information. Note that when applying the first category of attacks to craft the malicious model updates on fake clients, we assume that the attacker has access to the model updates of all genuine clients and the aggregation rule used by the server. In Section 8, we also study the scenarios where an attacker uses the global models to reconstruct synthetic data on the fake clients and uses the local models trained on them to perform attacks. We do not make these assumptions for Random, MPAF, and our PoisonedFL. Therefore, we give advantages to the first category of attacks.
@@
Fang [10]:  In this attack, the attacker crafts malicious model updates for fake clients such that the aggregated model update after attack deviates substantially from the before-attack one. Fang has different versions for different aggregation rules. We apply their Krum attack for Multi-Krum defense and TrMean attack for others.
@@
Opt. Fang [11]:  Following Fang, this attack also aims to maximize the deviation between the after-attack aggregated model update and before-attack one, but it uses a different way to solve the malicious model update. Specifically, a malicious model update is a variant of the average of all genuine clients’ model updates.
@@
A little is enough (LIE) [8]:  In LIE attack, fake clients craft their malicious model updates by adding a small amount of noise to the average of genuine model updates. Specifically, for each dimension $j$, LIE calculates the mean $\mu_{j}$ and standard deviation $\sigma_{j}$ among the genuine model updates, and the dimension $j$ of the malicious model update is set as $\mu_{j}+0.74\cdot\sigma_{j}$.
@@
Min-Max [11]:  In this attack, a fake client crafts its malicious model update such that its distance to any genuine model update is no larger than the maximum distance between any two genuine model updates.
@@
Min-Sum [11]:  Each fake client crafts its malicious model update such that the sum of the distances between the malicious model update and all genuine model updates is not larger than the sum of distances between any genuine model update and other genuine model updates.
@@
Random [9]:  In this attack, each fake client $i\in[n+1,n+m]$ sends a scaled random vector $\bm{g}_{i}^{t}=\lambda\cdot\bm{\epsilon}$ to the server, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\lambda$ is a scaling factor. Following prior work [9], we set $\lambda=1e6$ in our experiments.
@@
MPAF [9]:  In MPAF attack, during training round $t$, fake client $i\in[n+1,n+m]$ crafts its malicious model update as $\bm{g}_{i}^{t}=\lambda\cdot(\bm{w}^{\prime}-\bm{w}^{t-1}$), where $\bm{w}^{\prime}$ is an attacker-chosen random model, $\bm{w}^{t-1}$ is the global model in the previous training round $t-1$, and $\lambda$ is a scaling factor. Following prior work [9], we set $\lambda=1e6$ in our experiments.